HIGH-RESOLUTION INFRARED SPECTROSCOPY OF HCN-Mg_{n} CLUSTERS

\underline{PAUL L. STILES}, KLAAS NAUTA, DAVID T. MOORE and ROGER E. MILLER, Department of Chemistry, University of North Carolina, Chapel Hill, NC 27599.

Rotationally resolved infrared spectra have been obtained for the first time for metal complex-adsorbate systems. In particular, bare Mg clusters have been grown in helium nanodroplets, to which are added a single HCN molecule. The IR spectra associated with the C-H vibrational band have been recorded for HCN-Mg_{n} (n = 1 − 6). The results indicate that there is a qualitative change in the HCN-Mg cluster bonding (highly non-additive) in going from n = 2 to n = 4.